CH/Tech/96/01

Update

Report on Data Acquisition Tests

Author: Martin Freer

Date: 25 March 1996

In Canberra this January some further tests of the new data acquisition
system for Megha were performed. The main aim of the tests were to
evaluate the performance of the new data acquisition system as a whole
compared with the old/existing system. Due to difficulties in linking
the two systems together, so that data could be compared on an event by
event basis, the tests were not realised in their entirety. However, we
were able to make some important comparisons between the performance of
the two systems.
The tests followed a strip detector run, so we selected 8
strips, instrumenting the two ends, giving 16 channels of energy and
16 channels of time information. The signals from the preamplifiers
were split so that they could be simultaneously processed by the
two systems. We used pulser and elastic scattering data to perform
the measurements (30 MeV 12C beam off 12C and 197Au) targets.
The new data acquisition system contained 32 channels of David Smiths
programmable amplifiers, each containing a variety of shaping constants
(0.75, 1.1, 1.65, 2 us), this permitted an evaluation of the best
time constant for use with the position sensitive strip detectors.
The analogue signals were then processed by the Silena peak detect and
hold circuits followed by Dicks CAT (amplitude to time converter),
which converts the analogue signal to a representation which can be
processed by the Multi hit TDCs. One of the amplifiers also contained
a peak detect and hold circuit designed by David Smith.
Pulser Tests
------------
Matchsticks data provided by the BNC pulser were processed by the
new system to evaluate the linearity of the response. These measurements
indicate a extremely high degree of linearity over the dynamic range
1000-8000 (the TDCs have an 8k range) with the deviation from linear
being of the order of +/-2 channels (0.2% - in the worst case). This

deviation is significantly less that our energy resolution. The previous
tests had shown that the CAT circuit suffered from rather extreme
deviations from a linear response below channel 1000, with deviations
of up to 100. The current tests indicate that these deviations are now
much smaller, +/-8 channels in the worst case. This performance is now
comparable to, if not better than, that of the Silena ADC. There were
rather large offsets (250 channels) though, however these can be
adjusted and thus should not be considered to be a problem. There was
no apparent difference between the performance of the Silena and David
Smith peak detect circuits.
Elastic Scattering Data
-----------------------
Two types of tests were performed with the reaction products. The first
was a study of the complete response of the new system (which piped the
data words into the data stack in the old system), this was then
compared with the measurements with the old amplifiers and the Silena
ADCs. The second set of tests involved patching the new amplifiers into
Silena ADCs so that the data could be compared on an event by event
basis.
The spectra constructed from the first set of tests demonstrated that
there were no significant differences between the two systems, apart
from the higher thresholds in the new system. The energy resolutions
appeared to be approximately the same, 250 keV (new) 240 kev (old) for
the 30 MeV 12C ions. There appears to be a slight dependence of the
resolution  on the shaping time of the amplifier;
2.00 us - 250 keV
0.75 us - 235 keV
1.10 us - 220 keV
This difference is marginal and thus may not be significant, but does
suggest that the energy resolution is not compromised by the data
acquisition system.
The second set of tests allowed the same pulses processed by the
two systems to be compared on an event by event basis. Interestingly
there were some deviations between the two systems, this was particularly
for the amplifiers with longer shaping times (although not universally).
It is believed that this may by due to either the differing response
of the amplifiers to the noise or due to charge collection. Certainly
amplifiers with shaping constants closer to those of the old amplifiers
showed a much closer agreement. Importantly though, the energy resolution

is independent of the shaping time, and the position resolutions is
such that the positions calculated for the two sets of amps agree within
0.5 mm. Grids or masks would be required to provide a better determination
of the position resolution. These measurements were made with the same
shaping constants for the signals from either end of each strip. Hence,
although there exists some deviation between the amplifier responses,
there may be some cancelation of this effect in the calculation
of position and energy. It would be interesting to see the effect of
different time constants on either end of the strips.

In conclusion, I observe no significant differences between the two
systems. The energy and position resolution are similar. The performance
of the CAT circuit appears to be improved - particularly in terms of the
deviation at low channels, and there is no discernible difference between
the performance of the D. Smith and Silena peak detect circuits. Thus
from the point of view of the experimental measurements the new system
would appear to provide an equivalent response to that which we have
already. We were unable to compare the performance of the two systems
from the point of view of rate capacity, this of course is important,
but will have to wait until the new system is further developed.